# End-to-End Conformal Calibration for Optimization Under Uncertainty

### v1

@@ D.2 Data
We use the same dataset as Donti et al. [15] in our price forecasting for battery storage problem. In this dataset, the target $y\in\mathbb{R}^{24}$ is the hourly PJM day-ahead system energy price for 2011-2016, for a total of 2189 days. Unlike Donti et al. [15], though, we do not exclude any days whose electricity prices are too high (>500\$/MWh). Whereas Donti et al. [15] treated these days as outliers, our conditional robust optimization problem is designed to output robust decisions. For predicting target for a given day, the inputs $x\in\mathbb{R}^{101}$ include the previous day’s log-prices, the given day’s hourly load forecast, the previous day’s hourly temperature, the given day’s hourly temperature, and several calendar-based features such as whether the given day is a weekend or a US holiday.
@@
For the setting without distribution shift, we take a random 20% subset of the dataset as the test set; because the test set is selected randomly, it is considered exchangeable with the rest of the dataset. For the setting with distribution shift, we take the chronologically last 20% of the dataset as the test set; because load, electricity prices, and temperature all have distribution shifts over time, the test set is not exchangeable with the rest of the dataset. For each seed, we further use a 80/20 random split of the remaining data for training and calibration.
Portfolio optimization
@@
For the portfolio optimization task, we used synthetically generated data. We sample $x\in\mathbb{R}^{2},y\in\mathbb{R}^{2}$ from a mixture of three 4-D multivariate Gaussian distributions as used in [12]. Formally,
$\begin{bmatrix}x\\ y\end{bmatrix}\sim p_{a}\mathcal{N}(\mu_{a},\Sigma_{1})+p_{b}\mathcal{N}(\mu_{b},\Sigma_{b})+p_{c}\mathcal{N}(\mu_{c},\Sigma_{c})$
@@
$\Sigma_{a}$ $=\begin{bmatrix}1&0&0.37&0\\ 0&1.5&0&0\\ 0.37&0&2&0.73\\ 0&0&0.73&3\end{bmatrix},$ $\Sigma_{b}$ $=\alpha_{\text{GMM}}\Sigma_{a},$ $\Sigma_{c}$ $=\frac{1}{\alpha_{\text{GMM}}}\Sigma_{a}$
for some $\phi\in[0,1]$ and $\alpha_{\text{GMM}}\in[0,1]$. In our experiments, we used $\phi=0.7$ and $\alpha_{\text{GMM}}=0.9$ (it was unclear to us what values of $\phi$ and $\alpha_{\text{GMM}}$ were chosen by Chenreddy and Delage [12] for their experiments). For each random seed, we generate 2000 samples and use a (train, calibration, test) split of (600, 400, 1000).
D.3 Experimental results: Portfolio optimization
Figures 7 and 7 mirror Figures 3 and 3, except that the results here are on the portfolio problem with synthetic data. We again find that our E2E approach generally improves upon the ETO baselines at all uncertainty levels $\alpha$, with the exception of box uncertainty where all the methods achieve similar performance. Our PICNN-based uncertainty representation, when learned end-to-end, performs better than box uncertainty and comparably with ellipse uncertainty. The similarity in performance between E2E ellipsoidal uncertainty and E2E PICNN uncertainty is likely due to the underlying aleatoric uncertainty (i.e., the uncertainty in $\mathcal{P}(y\mid x)$) generally taking an ellipsoidal shape—the conditional distribution $\mathcal{P}(y\mid x)$ is still a Gaussian mixture model, and it tends to have a dominant mode. In terms of coverage, we find that all the models and training methodologies obtain coverage very close to the target level, confirming that the improvements in task loss performance from our E2E approach do not come at the cost of worse coverage.
Figure 6: Task loss performance (mean $\pm$1 stddev across 10 runs) for the portfolio optimization problem on synthetic data. Lower values are better.
Figure 7: Coverage (mean $\pm$1 stddev across 10 runs) for the portfolio optimization problem on synthetic data, with the dotted black line indicating the target coverage level $\alpha$.

### v2

@@ E.2 Data
We use the same dataset as donti_task-based_2017 in our price forecasting for battery storage problem. In this dataset, the target $y\in\mathbb{R}^{24}$ is the hourly PJM day-ahead system energy price for 2011-2016, for a total of 2189 days. Unlike donti_task-based_2017, though, we do not exclude any days whose electricity prices are too high (>500\$/MWh). Whereas donti_task-based_2017 treated these days as outliers, our conditional robust optimization problem is designed to output robust decisions. For predicting target for a given day, the inputs $x\in\mathbb{R}^{101}$ include the previous day’s log-prices, the given day’s hourly load forecast, the previous day’s hourly temperature, the given day’s hourly temperature, and several calendar-based features such as whether the given day is a weekend or a US holiday.
@@
For the setting without distribution shift, we take a random 20% subset of the dataset as the test set; because the test set is selected randomly, it is considered exchangeable with the rest of the dataset. For the setting with distribution shift, we take the chronologically last 20% of the dataset as the test set; because load, electricity prices, and temperature all have distribution shifts over time, the test set is not exchangeable with the rest of the dataset. Figure˜A4 illustrates the data from the distribution shift setting, where the electricity price tends to be lower and has lower variability in the test set compared to the training/validation splits. For each seed, we further use a 80/20 random split of the remaining data for training and calibration.
(a) Month of year vs. electricity price
(b) Hour of day vs. electricity price (shaded region shows $\pm$ 1 std dev)
(c) Daily high temperature (rounded to nearest 5F) vs. daily peak electricity price (shaded region shows $\pm$ 1 std dev)
Figure A4: Visualization of data from the battery storage problem under distribution shift. The train/val splits are sampled randomly from the range 2011-01-04 to 2015-10-20, whereas the test set comprises 2015-10-21 to 2016-12-31. These plots evidently show that the electricity price is generally lower in the test set and also has lower variability by both hour of day and the daily maximum temperature.
Portfolio optimization.
@@
For the portfolio optimization task, we used synthetically generated data. We sample $x\in\mathbb{R}^{2},y\in\mathbb{R}^{2}$ from a mixture of three 4-D multivariate Gaussian distributions as used in chenreddy_end--end_2024. Formally,
$\begin{bmatrix}x\\ y\end{bmatrix}\sim p_{a}\operatorname{\mathcal{N}}(\mu_{a},\Sigma_{a})+p_{b}\operatorname{\mathcal{N}}(\mu_{b},\Sigma_{b})+p_{c}\operatorname{\mathcal{N}}(\mu_{c},\Sigma_{c})$
@@
$\mu_{a}$ $=\mathbf{0}_{4},$ $\mu_{b}$ $=\begin{bmatrix}0&5&5&0\end{bmatrix}^{\top}$, $\mu_{c}$ $=\mu_{b},$
$\Sigma_{a}$ $=\begin{bmatrix}1&0&0.37&0\\ 0&1.5&0&0\\ 0.37&0&2&0.73\\ 0&0&0.73&3\end{bmatrix}$, $\Sigma_{b}$ $=\alpha_{\text{GMM}}\Sigma_{a},$ $\Sigma_{c}$ $=\frac{1}{\alpha_{\text{GMM}}}\Sigma_{a},$
for some $\phi\in[0,1]$ and $\alpha_{\text{GMM}}\in[0,1]$. In our experiments, we used $\phi=0.7$ and $\alpha_{\text{GMM}}=0.9$. (chenreddy_end--end_2024 do not disclose the values of $\phi$ and $\alpha_{\text{GMM}}$ chosen for their experiments.) For each random seed, we generate 2000 samples and use a (train, calibration, test) split of (600, 400, 1000).